The area of a square is decreased by 80%. By what percent did the perimeter decrease?
Answer by Alan3354(69443)

The area of a square is decreased by 80%. By what percent did the perimeter decrease?
--------
After decrease area = 0.2 of the original.
Area After/Area Before = 0.2
-----
Area = s^2
After s^2/Before s^2 = 0.2
After s/Before x = sqrt(0.2) =~ 0.447 = 44/7%
Side Decrease = 100 - 44.7 = 55.3%
Perimeter decrease = 55.3%